AgNO3 +CuCl2 --> AgCl + Cu(NO3)3

If 1.5l of 0.8M AgNO3 is added to 1.2l of 0.6M CuCl2 what is the concentration of AgCl in the new solution?


1
Expert's answer
2021-01-25T07:01:08-0500

C(AgCl)=C(AgNO3)=0.8 M

m= M×n=143.5×0.8=114.5 g

"\rho =114.5/(1.2+1.5)=42.5 g/L"